OpenStudy (anonymous):

let theta be angle in quadrant II such that sec(theta)=-10/7. Find the exact values of cot(theta) and sin(theta)

rvc (rvc):

1+tan^2theta=sec^2theta tantheta=1/cottheta

OpenStudy (anonymous):

I have \[\cot \theta=\frac{ 7 }{ \sqrt{51} }, \sin \theta=-\frac{ 51 }{ 10 }\]

OpenStudy (anonymous):

@rvc is that right?

rvc (rvc):

cot correct

OpenStudy (anonymous):

\[\sin \theta should be -\frac{ \sqrt{51} }{ 10 }\]

rvc (rvc):

cot is minus and sin is correct

rvc (rvc):

good!!